In a brood of 2000 cats, 5 have feline leukemia and the number of cats with leukemia doubles every day.
a. Write the number of cats with leukemia as a function of time (p=po(1+r)^t/n)
b. Graph the function over the first 30 days and show the window.
c. How many cats will have feline leukemia after 17 days?
d. On what day will the entire brood have feline leukemia?

Answer by stanbon(75887)

In a brood of 2000 cats, 5 have feline leukemia and the number of cats with leukemia doubles every day.
a. Write the number of cats with leukemia as a function of time (p=po(1+r)^t/n)
p(t) = 5*2^t
------
b. Graph the function over the first 30 days and show the window.
p(30) = 5*2^30
c. How many cats will have feline leukemia after 17 days?
p(30) = 5*2^17
d. On what day will the entire brood have feline leukemia?
Solve 5*2^t = 2000
---
2^t = 400
t = log(400)/log(2) = 8.64
Ans:: On the 9th day
---------------
Cheers,
Stan H.
----------